# Estimating stationary mass, frequency by frequency

### v2

@@ Abstract
Suppose we observe a trajectory of length $n$ from an $\alpha$-mixing stochastic process over a finite but potentially large state space. We consider the problem of estimating the probability mass placed by the stationary distribution of any such process on elements that occur with a certain frequency in the observed sequence. We estimate this vector of probabilities in total variation distance, showing universal consistency in $n$ and recovering known results for i.i.d. sequences as special cases. Our proposed methodology carefully combines the plug-in (or empirical) estimator with a recently-proposed modification of the Good–Turing estimator called WingIt, which was originally developed for Markovian sequences. En route to controlling the error of our estimator, we develop new performance bounds on WingIt and the plug-in estimator for $\alpha$-mixing stochastic processes. Importantly, the extensively used method of Poissonization can no longer be applied in our non i.i.d. setting, and so we develop complementary tools—including concentration inequalities for a natural self-normalized statistic of mixing sequences—that may prove independently useful in the design and analysis of estimators for related problems.
@@ 1 Introduction
Estimating the stationary distribution of a stochastic process from an $n$-length trajectory is a foundational problem in statistics and machine learning, with broad implications for fields such as ecology (Fisher et al., 1943), genomics (Favaro et al., 2012; Lijoi et al., 2007) and natural language processing (Church and Gale, 1991; Chen and Goodman, 1999; Ney et al., 1994). Historically, most research has focused on settings where the data is either i.i.d. or exchangeable, where classical techniques were based on so-called “add-constant” estimators (Laplace, 1814; Krichevsky and Trofimov, 1981). Other estimators that see appealing empirical performance are absolute discounting (Ney et al., 1994), Jelinek–Mercer smoothing (Jelinek, 1985), and the Good–Turing estimator (Good, 1953). These estimators proceed by first estimating the probability mass [^1] Note that $M^{\pi}_{\zeta}$ is a random functional of the stochastic process, since the set of symbols occurring $\zeta$ times in the observed sequence is random. $M^{\pi}_{\zeta}$ placed by the distribution $\pi$ on symbols occurring with each frequency $\zeta=0,1,\ldots,n$ in the sequence. One can then transform the estimate of the vector $(M^{\pi}_{\zeta})_{\zeta=0}^{n}$ to an estimator of $\pi$ itself; see Orlitsky and Suresh (2015). Among the aforementioned estimators, the Good–Turing estimator has been thoroughly analyzed in the i.i.d. setting, starting from the pioneering work of McAllester and Schapire (2000). In particular, these authors showed a universal result of the following form: For any i.i.d. sequence from a distribution $\pi$ defined on a finite alphabet, the Good–Turing estimator consistently estimates the probability mass $M^{\pi}_{\zeta}$ placed by the distribution on symbols occurring with each fixed frequency $\zeta$ in the sequence. Subsequent work (Drukh and Mansour, 2005; Acharya et al., 2013; Orlitsky and Suresh, 2015) refined these bounds and proposed a hybrid estimator, i.e. the Good–Turing estimator to approximate $M^{\pi}_{\zeta}$ for small $\zeta$ (i.e., to estimate the probability mass placed on small-frequency symbols) and the plug-in or empirical estimator to approximate $M^{\pi}_{\zeta}$ for large $\zeta$ (i.e., to estimate the probability mass placed on large-frequency symbols). In particular, the problem of estimating the probability vector $(M^{\pi}_{\zeta})^{n}_{\zeta=0}$ has been studied in various divergences including total variation, Kullback–Leibler, and chi-square. In addition to upper bounds, we also have an intricate understanding of the sharpness properties of various estimators as well as minimax lower bounds (Acharya et al., 2013; Orlitsky and Suresh, 2015). There has also been a significant body of follow-up work in the i.i.d. setting (see, e.g., Ohannessian and Dahleh, 2012; Painsky, 2023; Hao and Orlitsky, 2019).
@@
While the aforementioned papers have provided a detailed understanding of the i.i.d. setting, many real-world stochastic processes exhibit temporal dependencies. Examples abound in natural language processing, where sequences of words are sometimes better captured through hidden Markov modeling, and genomics, where DNA sequences are formed from interdependent base pairs. These dependencies in the stochastic process pose significant challenges for estimation of the stationary distribution. Indeed, the naive application of estimators such as Good–Turing is no longer valid even when the stochastic process is Markovian (Chandra et al., 2024) and different algorithmic tools are required for addressing temporal dependence (Hao et al., 2018; Skorski, 2020; Chandra et al., 2024). Motivated by this issue, recent work by Pananjady et al. (2024) studied the Markovian setting and introduced a “leave-a-window-out” variant of the Good–Turing estimator for estimating the stationary missing mass, i.e., the probability mass placed by the stationary distribution on elements that do not occur in the observed sequence. They showed that this variant of Good–Turing — which they termed windowed Good–Turing or WingIt — estimates the stationary missing mass with minimax optimal mean-squared error. Some extensions were also provided for estimation of the “small-count” probability, i.e., the mass placed by the distribution on elements occurring at most a certain number of times in the observed sequence. However, these results do not imply a satisfactory bound on distribution estimation. In particular, we still do not have an estimator for the stationary distribution in a frequency-by-frequency sense that estimates the vector of probabilities $(M^{\pi}_{\zeta})_{\zeta=0}^{n}$ consistently in any natural divergence measure.
@@
In this paper, we propose and theoretically analyze such an estimator. Specifically, we consider a stochastic process $X^{n}=(X_{1},X_{2},\ldots,X_{n})$ defined over a finite but unknown state space $\mathcal{X}$ in which we accommodate the regime $|\mathcal{X}|\gg n$. We assume that the process is $\alpha$-mixing —which subsumes mixing Markov chains and some hidden Markov models; see Section 2.1 to follow — and aim to estimate its stationary distribution $\pi$ in the aforementioned sense. Specifically, letting $M^{\pi}_{\zeta}$ denote the probability mass placed by the stationary distribution on elements occurring $\zeta$ times in $X^{n}$, our goal is to estimate the vector $(M^{\pi}_{\zeta})_{\zeta=0}^{n}$ in total variation distance. Importantly, we would like to develop a universal result like in the i.i.d. case, which shows that estimation can always be performed consistently in $n$. We show that such an estimator then naturally yields an estimator for the stationary distribution $\pi$ defined over the sample space $\mathcal{X}$ (see Lemma 1).
@@
  • In Theorem 1, we provide a risk bound on the proposed hybrid estimator, showing that with the appropriate parameter settings it attains total variation risk $\widetilde{\mathcal{O}}\left(n^{-1/6}\sqrt{\mathsf{t_{mix}}}\right)$ for any exponential $\alpha$-mixing process with mixing time $\mathsf{t_{mix}}$ (see Eqs. (1a) and (2)). This rate recovers known guarantees in the i.i.d. setting as a special case.
Our main result, Theorem 1, is proved through two key technical pieces of possible independent interest. First, Proposition 1, presented in Section 4, provides a high-probability bound on the $\ell_{1}$ error of the plug-in estimator. Notably, these results recover the known guarantees for the i.i.d. setting as special cases but our analysis does not rely on Poissonization, a common technique tailored to the i.i.d. setting used to handle dependent multiplicities of symbols. Instead, we develop an alternative approach based on blocking arguments, which decomposes the sequence into approximately independent blocks. In the process, we prove concentration inequalities for a natural self-normalized statistic for mixing stochastic processes.
Second, we establish a bound on the $\ell_{1}$-risk of the WingIt estimator for a general $\alpha$-mixing process in Proposition 2, presented in Section 4. Importantly, simply applying the worst-case bounds provided in Pananjady et al. (2024) would not yield Theorem 1, and so we provide a sharper bound on the $\ell_{1}$-risk that adapts to the behavior of the stationary distribution $\pi$. Specifically, if $\pi$ assigns low probability to low-frequency symbols, the bound naturally adjusts to reflect this, leading to a smaller error for the WingIt estimator.
The rest of this paper is organized as follows. Section 2 describes basic background and formalizes our problem statement. Section 3 describes the methodology for our estimator, and our main results for the analysis of this estimator are described in Section 4. In Section 5, we provide proofs of our main results, deferring statements and proofs of auxiliary lemmas to the appendix.
@@ 2.1 Model for the stochastic process
    Random duplication model: Suppose i.i.d. samples from $\pi$ are input to a random duplication model (Chandra and Thangaraj, 2024), which for each input $x$, outputs the duplicated sequence $(\underbrace{x,\ldots,x}_{k\text{ times}})$ with probability $\alpha$, and $x$ otherwise. The output process is ergodic and satisfies exponential $\alpha$-mixing, and the mixing time satisfies $\mathsf{t_{mix}}(\epsilon)=k$ for all $\epsilon>0$.
@@ 2.2 Frequency-by-frequency estimation of the stationary distribution
as the mass placed by $\pi$ on elements occurring exactly $\zeta$ times in the observed sequence $X^{n}$. Note that each scalar $M^{\pi}_{\zeta}(X^{n})$ is a random functional, since it depends both on the underlying distribution as well as on the realized sequence $X^{n}$. Denote the vector of count probabilities by
@@
While the problem of estimating the vector $M^{\pi}$ is interesting in itself, it is worth pausing to ask in what sense this yields an estimator for the stationary distribution $\pi$. To describe this, we recall the notion of a natural estimator due to Orlitsky and Suresh (2015), which is an estimator of $\pi$ that assign the same probability to all symbols appearing with the same frequency in $X^{n}$. Formally, define the set of all natural estimators
@@
Note that natural estimators are indeed intuitive, since if we know only that a sequence is $\alpha$-mixing and therefore reaches stationarity, then it is natural to assign the same probability to symbols that occur an equal number of times. In particular, any estimator $\widehat{M}:\mathcal{X}^{n}\to\Delta(\{0,1,\ldots,n\})$ of the count probability vector can be used to generate an estimator $\widehat{q}$ for the stationary distribution by dividing the mass $\widehat{M}_{\zeta}$ equally among all elements of $\mathcal{X}$ that occur $\zeta$ times in $X^{n}$. The estimator $\widehat{q}$ is natural by definition. The following lemma shows that this estimator is competitive with respect to the class of natural estimators $\mathcal{Q}^{\mathsf{nat}}$.
@@ 4 Main results
uniformly over all $\alpha$-mixing stochastic processes satisfying Eq. (1b).
@@
As a side remark, we note that Acharya et al. (2013) and Orlitsky and Suresh (2015) also analyze the error of their estimator in KL-divergence, and show that it achieves a rate of $\mathcal{O}(n^{-1/3})$ for IID sequences. By applying the reverse Pinsker inequality (Sason, 2015, Theorem 1) to Theorem 1 above, one can show that a minor modification of our estimator [^3] To apply the reverse Pinsker inequality, we need to ensure that our estimator does not assign probability equal to 00 to any $\zeta$. This can be done by adding $1$ to the total count of each $\zeta$, ensuring that $\widehat{M}_{\zeta}(\tau;\overline{\zeta})>1/n$ for all $\zeta$. attains KL risk $\widetilde{\mathcal{O}}((\tau^{3}/n)^{1/6})$ for general mixing sequences. While consistent, this rate is suboptimal, and proving a faster rate in KL error for our estimator remains an open problem.
As mentioned above, the proof of our main result relies on a concentration inequality for a self-normalized statistic of $\alpha$-mixing sequences, which we state below along with a more standard Bernstein-type inequality. We prove the lemma in Section 5.3.
@@
On the one hand, it is worth comparing the Bernstein bound in Lemma 2(a) to the Bernstein bound for $\alpha$-mixing sequences given in Merlevède et al. (2009, Theorem 2). The authors in that paper show that if $U^{n}$ is $\alpha$-mixing and satisfies $\alpha(\tau)\leq\exp(-2c\tau)$, then defining
@@
Such a self-normalized concentration inequality for exponential $\alpha$-mixing sequences may be of independent interest.

### v3

@@ Abstract
Suppose we observe a trajectory of length $n$ from an exponentially $\alpha$-mixing stochastic process over a finite but potentially large state space. We consider the problem of estimating the probability mass placed by the stationary distribution of any such process on elements that occur with a certain frequency in the observed sequence. We estimate this vector of probabilities in total variation distance, showing universal consistency in $n$ and recovering known results for i.i.d. sequences as special cases. Our proposed methodology—implementable in linear time—carefully combines the plug-in (or empirical) estimator with a recently-proposed modification of the Good–Turing estimator called WingIt, which was originally developed for Markovian sequences. En route to controlling the error of our estimator, we develop new performance bounds on WingIt and the plug-in estimator for exponentially $\alpha$-mixing stochastic processes. Importantly, the extensively used method of Poissonization can no longer be applied in our non i.i.d. setting, and so we develop complementary tools—including concentration inequalities for a natural self-normalized statistic of mixing sequences—that may prove independently useful in the design and analysis of estimators for related problems. Simulation studies corroborate our theoretical findings.
@@ 1 Introduction
Estimating the stationary distribution of a stochastic process from an $n$-length trajectory is a foundational problem in statistics and machine learning, with broad implications for fields such as ecology (Fisher et al., 1943), genomics (Favaro et al., 2012; Lijoi et al., 2007) and natural language processing (Church and Gale, 1991; Chen and Goodman, 1999; Ney et al., 1994). Historically, most research has focused on settings where the data is either i.i.d. or exchangeable, where classical techniques were based on so-called “add-constant” estimators (Laplace, 1814; Krichevsky and Trofimov, 1981). Other estimators that see appealing empirical performance are absolute discounting (Ney et al., 1994), Jelinek–Mercer smoothing (Jelinek, 1985), and the Good–Turing estimator (Good, 1953). These estimators proceed by first estimating the probability mass [^1] Note that $M^{\pi}_{\zeta}$ is a random functional of the stochastic process, since the set of symbols occurring $\zeta$ times in the observed sequence is random. $M^{\pi}_{\zeta}$ placed by the distribution $\pi$ on symbols occurring with each frequency $\zeta=0,1,\ldots,n$ in the sequence, and this estimation problem will form the focus of the current paper. One can then transform the estimate of the vector $(M^{\pi}_{\zeta})_{\zeta=0}^{n}$ to an estimator of $\pi$ itself; see Orlitsky and Suresh (2015). Among the aforementioned estimators, the Good–Turing estimator has been thoroughly analyzed in the i.i.d. setting, starting from the pioneering work of McAllester and Schapire (2000). In particular, these authors showed a universal result of the following form: For any i.i.d. sequence from a distribution $\pi$ defined on a finite alphabet, the Good–Turing estimator consistently estimates the probability mass $M^{\pi}_{\zeta}$ placed by the distribution on symbols occurring with each fixed frequency $\zeta$ in the sequence. Subsequent work (Drukh and Mansour, 2005; Acharya et al., 2013; Orlitsky and Suresh, 2015) refined these bounds and proposed a hybrid estimator, i.e. the Good–Turing estimator to approximate $M^{\pi}_{\zeta}$ for small $\zeta$ (i.e., to estimate the probability mass placed on small-frequency symbols) and the plug-in or empirical estimator to approximate $M^{\pi}_{\zeta}$ for large $\zeta$ (i.e., to estimate the probability mass placed on large-frequency symbols). In addition, the problem of estimating the missing mass, $M^{\pi}_{0}$, has garnered significant interest in the i.i.d. setting and has been analyzed thoroughly (McAllester and Ortiz, 2003; Berend and Kontorovich, 2012, 2013). In particular, the problem of estimating the probability vector $(M^{\pi}_{\zeta})^{n}_{\zeta=0}$ has been studied in various divergences including total variation, Kullback–Leibler, and chi-square. In addition to upper bounds, we also have an intricate understanding of the sharpness properties of various estimators as well as minimax lower bounds (Acharya et al., 2013; Orlitsky and Suresh, 2015). There has also been a significant body of follow-up work in the i.i.d. setting (see, e.g., Ohannessian and Dahleh, 2012; Painsky, 2023; Hao and Orlitsky, 2019).
@@
While the aforementioned papers have provided a detailed understanding of the i.i.d. setting, many real-world stochastic processes exhibit temporal dependencies. Examples abound in natural language processing, where sequences of words are sometimes better captured through hidden Markov modeling, and genomics, where DNA sequences are formed from interdependent base pairs. These dependencies in the stochastic process pose significant challenges for estimation of the stationary distribution. Indeed, the naive application of estimators such as Good–Turing is no longer valid even when the stochastic process is Markovian (Chandra et al., 2024) and different algorithmic tools are required for addressing temporal dependence (Hao et al., 2018; Skorski, 2020; Chandra et al., 2024). Motivated by this issue, recent work by Pananjady et al. (2024) studied the Markovian setting and introduced a “leave-a-window-out” variant of the Good–Turing estimator for estimating the stationary missing mass, i.e., the probability mass placed by the stationary distribution on elements that do not occur in the observed sequence. They showed that this variant of Good–Turing — which they termed windowed Good–Turing or WingIt — estimates the stationary missing mass with minimax optimal mean-squared error. Some extensions were also provided for estimation of the “small-count” probability, i.e., the mass placed by the distribution on elements occurring at most a certain number of times in the observed sequence. However, these results do not imply a satisfactory bound on distribution estimation. In particular, we still do not have an estimator for the stationary distribution in a frequency-by-frequency sense that estimates the vector of probabilities $(M^{\pi}_{\zeta})_{\zeta=0}^{n}$ consistently in any natural divergence measure. Finally, we mention that besides frequency-by-frequency stationary distribution estimation, researchers have studied various learning problems involving other parameters of Markov chains (e.g., Hao et al., 2018; Han et al., 2023; Wolfer and Kontorovich, 2019).
@@
In this paper, we propose and theoretically analyze an estimator for the stationary distribution in a frequency-by-frequency sense. Specifically, we consider a stochastic process $X^{n}=(X_{1},X_{2},\ldots,X_{n})$ defined over a finite but unknown state space $\mathcal{X}$ in which we accommodate the regime $|\mathcal{X}|\gg n$. We assume that the process is exponentially $\alpha$-mixing —which subsumes mixing Markov chains and some hidden Markov models; see Section 2.1 to follow — and aim to estimate its stationary distribution $\pi$ in the aforementioned sense. Specifically, letting $M^{\pi}_{\zeta}$ denote the probability mass placed by the stationary distribution on elements occurring $\zeta$ times in $X^{n}$, our goal is to estimate the vector $(M^{\pi}_{\zeta})_{\zeta=0}^{n}$ in total variation distance. Importantly, we would like to develop a universal result like in the i.i.d. case, which shows that estimation can always be performed consistently in $n$. We show that such an estimator then naturally yields an estimator for the stationary distribution $\pi$ defined over the sample space $\mathcal{X}$ (see Lemma 1).
@@
  • In Theorem 1, we provide a risk bound on the proposed hybrid estimator, showing that with the appropriate parameter settings it attains total variation risk $\widetilde{\mathcal{O}}\left(n^{-1/6}\sqrt{\mathsf{t_{mix}}}\right)$ for any exponentially $\alpha$-mixing process with mixing time $\mathsf{t_{mix}}$ (see Eqs. (1a) and (2)). This rate recovers known guarantees in the i.i.d. setting as a special case.
Our main result, Theorem 1, is proved through two key technical pieces of possible independent interest. First, Proposition 1, presented in Section 4, provides a high-probability bound on the $\ell_{1}$ error of the plug-in estimator. Notably, these results recover the known guarantees for the i.i.d. setting as special cases, but our analysis does not rely on Poissonization, a common technique tailored to the i.i.d. setting used to handle dependent multiplicities of symbols. Instead, we develop an alternative approach based on blocking arguments, which decomposes the sequence into approximately independent blocks. In the process, we prove concentration inequalities for a natural self-normalized statistic for mixing stochastic processes.
Second, we establish a bound on the $\ell_{1}$-risk of the WingIt estimator for a general exponentially $\alpha$-mixing process in Proposition 2, presented in Section 4. Importantly, simply applying the worst-case bounds provided in Pananjady et al. (2024) would not yield Theorem 1, and so we provide a sharper bound on the $\ell_{1}$-risk that adapts to the behavior of the stationary distribution $\pi$. Specifically, if $\pi$ assigns low probability to low-frequency symbols, the bound naturally adjusts to reflect this, leading to a smaller error for the WingIt estimator.
The rest of this paper is organized as follows. Section 2 describes basic background and formalizes our problem statement. Section 3 describes the methodology for our estimator, and our main results for the analysis of this estimator are described in Section 4. In Section 5, we simulate our estimator and compare its performance to alternatives that are available in the literature. In Section 6, we provide proofs of our main results, deferring statements and proofs of auxiliary lemmas to the appendix.
@@ 2.1 Model for the stochastic process
    Random duplication model: Suppose i.i.d. samples from $\pi$ are input to a random duplication model (Chandra and Thangaraj, 2024), which for each input $x$, outputs the duplicated sequence $(\underbrace{x,\ldots,x}_{k\text{ times}})$ with probability $\alpha$, and $x$ otherwise. The output process is ergodic and satisfies exponentially $\alpha$-mixing, and the mixing time satisfies $\mathsf{t_{mix}}(\epsilon)=k$ for all $\epsilon>0$.
@@ 2.2 Frequency-by-frequency estimation of the stationary distribution
as the mass placed by $\pi$ on elements occurring exactly $\zeta$ times in the observed sequence $X^{n}$. Note that each scalar $M^{\pi}_{\zeta}(X^{n})$ is a random functional, since it depends both on the underlying distribution and on the realized sequence $X^{n}$. Denote the vector of count probabilities by
@@
While the problem of estimating the vector $M^{\pi}$ is interesting in itself, it is worth pausing to ask in what sense this yields an estimator for the stationary distribution $\pi$. To describe this, we recall the notion of a natural estimator due to Orlitsky and Suresh (2015), which is an estimator of $\pi$ that assigns the same probability to all symbols appearing with the same frequency in $X^{n}$. Formally, define the set of all natural estimators
@@
Note that natural estimators are indeed intuitive, since if we know only that a sequence is exponentially $\alpha$-mixing and therefore reaches stationarity, then it is natural to assign the same probability to symbols that occur an equal number of times. In particular, any estimator $\widehat{M}:\mathcal{X}^{n}\to\Delta(\{0,1,\ldots,n\})$ of the count probability vector can be used to generate an estimator $\widehat{q}$ for the stationary distribution by dividing the mass $\widehat{M}_{\zeta}$ equally among all elements of $\mathcal{X}$ that occur $\zeta$ times in $X^{n}$. The estimator $\widehat{q}$ is natural by definition. The following lemma shows that this estimator is competitive with respect to the class of natural estimators $\mathcal{Q}^{\mathsf{nat}}$.
@@ 3.1 WingIt Estimator
We now present an explicit implementation of the WingIt estimator defined in Eq. (9). We implement the WingIt estimator in parallel for all values of $\zeta\in\{0,\ldots,n\}$ and return the $n+1$-dimensional vector $\widehat{M}_{\textsc{WingIt}}$. This implementation runs in $\mathcal{O}(n\tau)$ time, where $\tau$ is the window size used. Given a sequence $X^{n}$ and a natural number $\tau$, our implementation proceeds via two passes through the data as shown in Algorithm 1.
Algorithm 1 Linear time implementation of the WingIt estimator.
0:  Sequence $X^{n}=(X_{1},\ldots,X_{n})$, Natural number $\tau$ (window size)
1:  Initialize locations as a dictionary
2:  for $i=1,\ldots,n$ do
3:     if $X_{i}\notin$ locations then
4:        Initialize locations$[X_{i}]$ as a list
5:     end if
6:     Append $i$ to locations$[X_{i}]$
7:  end for
8:  $\widehat{M}_{\textsc{WingIt}}\leftarrow[\underbrace{0,\ldots,0}_{n+1}]$
9:  for $i=1,\ldots,n$ do
10:     count $\leftarrow 0$
11:     for $j=\max(0,i-\tau),\ldots,\min(i+\tau,n)$ do
12:        if $X_{j}==X_{i}$ then
13:           count $\leftarrow$ count $+1$
14:        end if
15:     end for
16:     $\zeta\leftarrow\texttt{len}(\texttt{locations}[X_{i}])-$ count
17:     $\widehat{M}_{\textsc{WingIt}}[\zeta]\leftarrow\widehat{M}_{\textsc{WingIt}}[\zeta]+1/n$
18:  end for
19:  return  $\widehat{M}_{\textsc{WingIt}}$
The first loop in Algorithm 1 (Steps 2-7) requires a single pass through the data and thus runs in $\mathcal{O}(n)$ time. This loop stores the locations of each symbol in the dictionary locations. In Step 8 we initialize the vector of count probabilities $\widehat{M}_{\textsc{WingIt}}$ as a vector of zeros of size $n+1$ corresponding to each frequency. The $\zeta$-th element of this vector, denoted by $\widehat{M}_{\textsc{WingIt}}[\zeta]$, corresponds to the quantity $\widehat{M}_{\textsc{WingIt},\zeta}(\tau)$ defined in Eq. (9). The second loop in this algorithm (Steps 9-18) runs in $\mathcal{O}(n\tau)$ time. The outer loop performs a single pass through the data, which takes $\mathcal{O}(n)$ time, and the inner loop (Steps 11-15) counts the number of occurrences of a symbol in a window around its current occurrence, which takes $\mathcal{O}(\tau)$ time. The variable $\zeta$ counts the number of occurrences of the symbol outside the window in Step 16, which can be implemented in $\mathcal{O}(1)$ time on Python. Finally, the algorithm returns the vector of count probabilities of the WingIt estimator and runs in $\mathcal{O}(n\tau)$ time.
@@ 3.2 Plug-in estimator
Clearly, the plug-in estimator (simultaneously for all values of $\zeta$) is linear-time implementable.
@@ 3.3 Combined estimator
The combined estimator can be implemented in linear time using the linear time implementation of the plug-in estimator and that of the WingIt estimator in Algorithm 1.
@@ 4 Main results
uniformly over all exponentially $\alpha$-mixing stochastic processes satisfying Eq. (1b).
@@
As a side remark, we note that Acharya et al. (2013) and Orlitsky and Suresh (2015) also analyze the error of their estimator in KL-divergence, and show that it achieves a rate of $\mathcal{O}(n^{-1/3})$ for IID sequences. By applying the reverse Pinsker inequality (Sason, 2015, Theorem 1) to Theorem 1 above, one can show that a minor modification of our estimator [^3] To apply the reverse Pinsker inequality, we need to ensure that our estimator does not assign 00 probability to any $\zeta$. This can be done by adding $1$ to the total count of each $\zeta$, ensuring that $\widehat{M}_{\zeta}(\tau;\overline{\zeta})>1/n$ for all $\zeta$. attains KL risk $\widetilde{\mathcal{O}}((\tau^{3}/n)^{1/6})$ for general mixing sequences. While consistent, this rate is suboptimal, and proving a faster rate in KL error for our estimator remains an open problem.
As mentioned above, the proof of our main result relies on a concentration inequality for a self-normalized statistic of exponentially $\alpha$-mixing sequences, which we state below along with a more standard Bernstein-type inequality. We prove the lemma in Section 6.3.
@@
On the one hand, it is worth comparing the Bernstein bound in Lemma 2(a) to the Bernstein bound for exponentially $\alpha$-mixing sequences given in Merlevède et al. (2009, Theorem 2). The authors in that paper show that if $U^{n}$ is $\alpha$-mixing and satisfies $\alpha(\tau)\leq\exp(-2c\tau)$, then defining
@@
Such a self-normalized concentration inequality for exponentially $\alpha$-mixing sequences may be of independent interest.